OpenStudy (anonymous):

Homework Help - Thanks! Differentiate: a) g(x)=2e^(-x/2) b) h(x)=In(x^2+3x)

OpenStudy (bibby):

b looks like simple chain rule a looks like u substitution

OpenStudy (anonymous):

Thank you!

OpenStudy (bibby):

would you like to actually work through them?

OpenStudy (anonymous):

Yes please

OpenStudy (bibby):

We'll start with b. recall that the derivative of ln(x) = 1/x As such, the derivative of In(x^2+3x) = 1/(x^2+3x) but that's not enough. the chain rule applies here so we multiply by the derivative of (x^2+3x)

OpenStudy (anonymous):

I am writing this down, its been 20 years since I have been in school so I am slowly learning again. I just need to get through this last math class. thanks so much...

OpenStudy (bibby):

Ah, well this should be a decent enough place for actually learning the math. What you want to do is get the elementary derivatives/formulas memorized. Keep them up in a tab while you work through problems. 2e^(-x/2) things to note: * the 2 is a constant multiplier. the derivative of 2g(x) = 2g'(x) [it just means we can ignore it and leave it out while we take the derivative of the rest] * we know the derivative of e^x but not e^(-x/2)

OpenStudy (anonymous):

Good advice.

OpenStudy (bibby):

2e^(-x/2) if we make u = -x/2 du = -1dx so we have to differentiate 2e^(u). the derivative of e^x is e^x. 2e^u * du because we have to multiply by the derivative of u.

OpenStudy (anonymous):

ok...you have been most helpful!

OpenStudy (bibby):

http://www.wolframalpha.com/input/?i=derive+2e%5E%28-x%2F2%29++ anyhow I have to go to school. check this out and feel free to message me if you have things you need help with

OpenStudy (anonymous):

Will do thanks!